OpenStudy (anonymous):

Can some one help me solve 0.319 = (10.7 + K ) over 46 im having trouble. Can some please help try to solve it ??

OpenStudy (anonymous):

0.319 =(10.7 + K)/46 first multiply the equation by 46 (0.319)(46) = (46)(10.7 + K)/46 14.674 = 10.7 + K can you take it from here ?

OpenStudy (anonymous):

all you do now is subtract 10.7 from each side and you have found K.

OpenStudy (anonymous):

Thank you so much you really helped me im going to give a medal if I can lol

OpenStudy (anonymous):

glad I could help :)

OpenStudy (anonymous):

:)